OpenStudy (anonymous):

Determine the magnitude of the force between an electric dipole with a dipole moment of 9e-29 C m and an electron. The electron is positioned r = 16 nm from the centre of the dipole, along the dipole axis. Hint: Assume that r >> d, with d the charge separation distance in the dipole.

OpenStudy (jamesj):

First of all, do you know how to calculate the electric field E, at this point where the electron is located? What you don't know is direction of the field--either towards or away from the dipole--but you should be able to calculate it. Once you know the magnitude of E there--call it E also for now--then the magnitude of the force on the electron is of course F = Ee, where e is our friend 1.6 x 10^-19

OpenStudy (anonymous):

ok i tried that but it didnt work :( and what about the r value, it has to be in there some how doesnt it?

OpenStudy (jamesj):

What is the expression for E, in general for a dipole?

OpenStudy (anonymous):

i figured it out, from a friend, its F=eE where E=9*10^9 * 2*P /r^3 where p=9*10^-29, r=16*10^-9 and r is our friend 1.6*10^-19 as you said :)

OpenStudy (anonymous):

thank you anyways :) i appreciate it